OpenStudy (anonymous):

Let n be an odd natural number. Prove that there exist polynomials pn(z),qn(z) such that for all real x, pn(cosx)=cos(nx) qn(sinx)=sin(nx) here pn and qn means p sub n & q sub n

OpenStudy (anonymous):

Notice \[\cos (3 x)=\cos ^3(x)-3 \sin ^2(x) \cos (x)=\\ \cos ^3(x)-3 \cos (x) \left(1-\cos ^2(x)\right)=\\4 \cos ^3(x)-3 \cos (x) \]

OpenStudy (anonymous):

The general case can be done by induction

OpenStudy (anonymous):

You will have to do the sin and cos separately

OpenStudy (anonymous):

You also can do it by using complex powers e \[\Large \cos(nx) =\frac { e^{i n x} +e^{-i n x} }{2}=\\ \Large \frac { \left(e^{i x} \right)^n+\left(e^{-i x} \right)^n}{2} \]

OpenStudy (anonymous):

\[\Large \cos(nx) =\frac { e^{i n x} +e^{-i n x} }{2}=\\ \Large \frac { \left(\cos( x) + i \sin( x) \right)^n+\left(\cos(x) -i \sin(x) \right)^n}{2}= \Large \]

OpenStudy (anonymous):

Use after the binomial Theorem See this for details http://mathworld.wolfram.com/Multiple-AngleFormulas.html

OpenStudy (perl):

are there are no conditions on p_n (cosx) such that it is an odd polynomial?

OpenStudy (anonymous):

Here is the case for n=11 \[\Large \cos (11 x)=1024 \cos ^{11}(x)-2816 \cos ^9(x)+\\ \Large 2816 \cos ^7(x)-1232 \cos ^5(x)+220 \cos ^3(x)-11 \cos (x) \]

OpenStudy (anonymous):

n=12 \[ \cos (12 x)=2048 \cos ^{12}(x)-6144 \cos ^{10}(x)+\\6912 \cos ^8(x)-3584 \cos ^6(x)+840 \cos ^4(x)-72 \cos ^2(x)+1 \]

OpenStudy (anonymous):

It will be a polynomial with degree n

OpenStudy (anonymous):

If you want to prove it by induction, use Chebyshev's recurrence formulae \[ \Large \cos (n x)=2 \cos (x) \cos ((n-1) x)-\cos ((n-2) x)\\ \Large \sin (n x)=2 \cos (x) \sin ((n-1) x)-\sin ((n-2) x)\\ \] @experimentX

OpenStudy (anonymous):

\[ \sin (11 x)=-1024 \sin ^{11}(x)+\\2816 \sin ^9(x)-2816 \sin ^7(x)+\\1232 \sin ^5(x)-220 \sin ^3(x)+11 \sin (x) \]

OpenStudy (anonymous):

I believe, you only need n odd for the sin part

OpenStudy (anonymous):

@satellite73